[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Apr 27, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Apr 27 15:36:54 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. Security Camera display in WFO MFL and TAFB were only showing half of  
the security cameras. Craig Mattocks opened a helpdesk ticket, and NCO  
resolved.
HSU:
1. Today is the last day of WFO storm surge testing, in which Mike Brennan  
submits a test psurge run and the affected WFOs verify ingest. Testing has  
been successful so far.

2. Matt Sardi discovered issues with the TILEFISH X server config file  
(xorg.conf), which could potentially cause N-AWIPS performance degradation  
similar to what TAFB experienced on BALLYHOO. A helpdesk ticket has been  
opened.
Other items:
1. Reminder, NCO is performing upgrades on the core Internet routers in  
College Park, MD today. So far, effects from Round 1 this morning have been  
minimal at NHC.

There will be a second period of potential impact from 1900 UTC (3pm ET) to  
2015 UTC (4:15 pm ET). Outages during these 75 minutes are not anticipated,  
but it is high risk work that they want users to be aware of.

2. NCO, Joe Zajic (HQ) and Monica Bozeman are still troubleshooting  
multiple issues with GOES-16 data flow. Expect sporadic data outages  
throughout the day for various channel numbers until this is fixed.
NHC RFCs:
None.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170427/95b23326/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list